Parts e, f, and both g please

Question 4: In the market for flowers. demand is given by: P = 40 - Q. The market supply is given by: P=. 1In this question, demand=p=40-1/400 supply=p=1/400 A) At equilibrium demand =supply so, 40-1/4Q=1/40Q: We get the value of Q=Now suppose the government would like to boost employment in the flower industry, so they implement a subsidy (or negative ta
Answer #1

(e)

After subsidy, new supply function is

P + 2 = (Q/40)

P = (Q/40) - 2

P = 0.025Q - 2

Equating with demand,

40 - (Q/10) = 0.025Q - 2

40 - 0.1Q = 0.025Q - 2

0.125Q = 42

Q = 336

(f)

P = 40 - 0.1 x 336 = 40 - 33.6 = 6.4 (Price paid by buyers)

(g)

Price received by sellers = 6.4 + 2 = 8.4

Share of subsidy for buyers = 8 - 6.4 = 1.6

Share of subsidy for sellers = 8.4 - 8 = 0.4

Since buyers' share of subsidy is higher, it means demand is more elastic than supply.

(g)

Cost of subsidy = 2 x 336 = 672

(h)

DWL = (1/2) x Subsidy x Change in Q = (1/2) x 2 x (336 - 320) = 1 x 16 = 16